OpenStudy (anonymous):

How many solutions does this system of equations have? x-3y=6 9y=3x-18

OpenStudy (anonymous):

??

OpenStudy (callisto):

rearrange the first equation x-3y=6 => x-3y-6=0 Consider the second equation 9y=3x-18 => 3x-9y-18=0 Take out 3 as the common factor and divide both sides by 3 3(x-3y-6) =0 x-3y-6 =0 So you'll get 2 same equations That means they are the same straight line, which have infinitely many solutions

OpenStudy (anonymous):

How many solutions does this system of equations have? x-4y=13 5x-20y=63